# GOLLuM: Gaussian Process Optimized LLMs – Reframing LLM Finetuning through Bayesian Optimization

### v2

@@ F.1 BO Initialization
This starting point contrasts with scenarios where good conditions are already known, in which case optimization reduces to local exploitation rather than global search. Reaching high-yielding conditions from poor initial data is substantially more challenging and better reflects real-world discovery pipelines. Compounding this difficulty is the negative bias in the scientific literature: failed or low-yield experiments are rarely reported, making published datasets inherently skewed toward successful outcomes 97.
@@ F.2 BO setup
Following the initialization, we run the BO loop for 50 iterations with batch size 1. We repeat each experiment configuration for 20 times with different seed values (1-20) to obtain robust performance metrics. Our choice for the GP kernel is Matérn-5/2, based on its demonstrated effectiveness on both continuous and discrete design spaces 37. For balancing the exploration and exploitation we employ the expected improvement acquisition function.
@@ F.3 Implementation Details
Surrogate Models.
We implement both fixed-feature and finetuned surrogate models as subclasses of SingleTaskGP from botorch 98. For fixed-feature GPs, the inputs are LLM embeddings, chemistry-related representations (fingerprints, DRFP) or default parameters (one-hot encoded categorical variables, numerical values), while we learn the GP kernel hyperparameters $\theta=\{\ell,\sigma^{2},\sigma_{n}^{2},c\}$ by maximizing the marginal likelihood using the L-BFGS-B optimizer provided by BoTorch’s fit_gpytorch_mll routine. We employ the Matérn-5/2 kernel 99 with initialization: $\ell=1.0$, $\sigma^{2}=1.0$, and $\sigma_{n}^{2}=1.0\times 10^{-4}$. The optimizer runs with multiple restarts, as part of BoTorch’s default behavior.
For deep kernel GPs, we extend the surrogate to jointly optimize both GP and LLM parameters. We build a custom DeepGP class that incorporates a finetuning model (PEFT adapter and/or projection head) jointly trained via AdamW 100. We optimize GP and LLM parameters using separate learning rates ($2\times 10^{-1}$ for GP, $2\times 10^{-3}$ for LLM) with a shared weight decay of $1\times 10^{-3}$. We apply gradient clipping with a max norm of 1.0 and decay the learning rates using a StepLR scheduler with a decay factor of 0.95.
PEFT Configuration.
@@
Tracking and Seeding.
All experiments are seeded using seed_everything from pytorch_lighting 101 with values 1-20. Acquisition optimization is deterministic over the candidate pool. We do not augment the data or use any stochastic featurization. We use wandb 102 for tracking experiments, managing seeds, logging losses, metrics, learning rates, images, and running configuration sweeps.

### v3

@@ F.1 BO Initialization
This starting point contrasts with scenarios where good conditions are already known, in which case optimization reduces to local exploitation rather than global search. Reaching high-yielding conditions from poor initial data is substantially more challenging and better reflects real-world discovery pipelines. Compounding this difficulty is the negative bias in the scientific literature: failed or low-yield experiments are rarely reported, making published datasets inherently skewed toward successful outcomes 105.
@@ F.2 BO setup
Following the initialization, we run the BO loop for 50 iterations with batch size 1. We repeat each experiment configuration for 20 times with different seed values (1-20) to obtain robust performance metrics. Our choice for the GP kernel is Matérn-5/2, based on its demonstrated effectiveness on both continuous and discrete design spaces 22. For balancing the exploration and exploitation we employ the expected improvement acquisition function.
@@ F.3 Implementation details
Surrogate Models
We implement both fixed-feature and finetuned surrogate models as subclasses of SingleTaskGP from botorch 106. For fixed-feature GPs, the inputs are LLM embeddings, chemistry-related representations (fingerprints, DRFP) or default parameters (one-hot encoded categorical variables, numerical values), while we learn the GP kernel hyperparameters $\theta=\{\ell,\sigma^{2},\sigma_{n}^{2},c\}$ by maximizing the marginal likelihood using the L-BFGS-B optimizer provided by BoTorch’s fit_gpytorch_mll routine. We employ the Matérn-5/2 kernel 107 with initialization: $\ell=1.0$, $\sigma^{2}=1.0$, and $\sigma_{n}^{2}=1.0\times 10^{-4}$. The optimizer runs with multiple restarts, as part of BoTorch’s default behavior.
For deep kernel GPs, we extend the surrogate to jointly optimize both GP and LLM parameters. We build a custom DeepGP class that incorporates a finetuning model (PEFT adapter and/or projection head) jointly trained via AdamW 108. We optimize GP and LLM parameters using separate learning rates ($2\times 10^{-1}$ for GP, $2\times 10^{-3}$ for LLM) with a shared weight decay of $1\times 10^{-3}$. We apply gradient clipping with a max norm of 1.0 and decay the learning rates using a StepLR scheduler with a decay factor of 0.95.
PEFT Configuration
@@
Tracking and seeding
All experiments are seeded using seed_everything from pytorch_lighting 109 with values 1-20. Acquisition optimization is deterministic over the candidate pool. We do not augment the data or use any stochastic featurization. We use wandb 110 for tracking experiments, managing seeds, logging losses, metrics, learning rates, images, and running configuration sweeps.